OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) + 6

OpenStudy (anonymous):

This video should help: http://www.youtube.com/watch?v=d3OYGEi6m-4

OpenStudy (anonymous):

3x+3=-2x+2+6 5x=5 X=5/5 X=1